OpenStudy (anonymous):

\[\int\frac{ dx }{ x^2+x^{3/2}-1 }\]

OpenStudy (anonymous):

int_{}^{}\frac{ dx }{ x^2+x^{3/2}-1 }\[\int\limits_{}^{}\frac{ dx }{ x^2+x^{3/2}-1 }\]

OpenStudy (anonymous):

@experimentX advice him take the perfect square of terms is the right one?

OpenStudy (experimentx):

huh ... I think x=t^2 subs

OpenStudy (experimentx):

wolf gives weird solution http://www.wolframalpha.com/input/?i=integrate+1%2F%28x%5E2++%2B+x%5E%283%2F2%29+-+1%29+ let me check on maple.

OpenStudy (experimentx):

maple says the same ... weirder than Wolf's

OpenStudy (anonymous):

LOL.. try my way. I remember that once my pro solve something like: x^3/2 = 2*1/2 x^2*x^1/2. so the third term of perfect square is 1/2x^1/2. and something more. I am not sure so I dare not to say restlessly, but if both math tools say something weird like that, why don't we try?

OpenStudy (experimentx):

are you trying to complete square?

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

i don't think it can be done ... if the second term were 'x' then it would have been easier. x^3/2 = 2*1/2 x^2*x^1/2 <--- should be x*x^(1/2) this would give out x^(1/4) at the end.

OpenStudy (anonymous):

yes, my bad. let me try again

OpenStudy (experimentx):

wolf decomposes it into partial fraction and integrates it where you get that log ... but i've learnt not to trust machines.

OpenStudy (anonymous):

LOL , sir, please, don't confuse me. you told me that you depend on machine too much, now i heard another thing, LOL... which one is correct?

OpenStudy (amistre64):

i asked the lawnmower to do the dishes once .... biiiig mistake

OpenStudy (anonymous):

@amistre64 what do you mean?

OpenStudy (amistre64):

experiment has learnted to not trust machines :)

OpenStudy (anonymous):

do you have any idea? I am waiting for help at my post. but stuck here with helper

OpenStudy (experimentx):

@Hoa you doing linear algebra bilinear forms?? @amistre64 I depended on wolf too much and sometimes I got busted ;(((

OpenStudy (anonymous):

not touch that part yet.

OpenStudy (amistre64):

im not seeing a "simple" solution yet. i know that sometimes, if you work the division you get a polymonial ... sometimes

OpenStudy (experimentx):

without target I am lost @cahit do you have answersheet?

OpenStudy (anonymous):

no i have not any answer or answer sheet